OpenStudy (anonymous):

Compute f '(a) algebraically for the given value of a. f(x) = x2 − 5; a = 3

OpenStudy (anonymous):

f'(x) = 2*x f'(3) = 2*3 = 6

OpenStudy (nurali):

f(x) = x2 − 5 Apply Derivative Formula f'(x)=2x f'(a)=2a a=3 f'(3)=2*3 f'(3)=6

OpenStudy (anonymous):

@Ashleyisakitty Do you understand how to differentiate (find the derivative)?

OpenStudy (anonymous):

\[ f'(a) = \lim_{x\to a}\frac{f(x)-f(a)}{x-a}= \lim_{x\to a}\frac{(x^2-5)-(a^2-5)}{x-a} \]

OpenStudy (anonymous):

\[ \lim_{x\to a}\frac{x^2-a^2}{x-a} = \lim_{x\to a}\frac{(x-a)(x+a)}{x-a}=\lim_{x\to a}x+a = a+a = 2a \]

OpenStudy (anonymous):

So since \[ f'(a) = 2a \]And \(a=3\), we get \[ f'(3) = 2(3)=6 \]Just as the others have said.